[Date Prev][Date Next][Thread Prev][Thread Next][Date Index][Thread Index]

Re: MK111c hair pulling problem



Fergus, I'm forwarding your message below to the group-- I feel sure 
someone here can help you.  I've also subscribed you to the mailing
list.  Good luck.

--- Forwarded mail from fmcarthy at indigo.ie (fergus mc carthy)

Hi Rob,
      I hope you can help me with this problem i have with our 14 year old
machine.The problem is as follows,when you set up on a open gate your pec
levels(ie) just below clip level as if you were checking shading alignment
and  stored that scene in a frame store then set up a split screen so as to
compere the reference image with the live image.At this point levels begin
to change either up or down on all three channels,this change is very
subtle and could take 5 mins to change or 5 hours.As you will appreciate
this makes grading impossible.In my attempt to sort this problem to date i
have removed the grading system renaissance from the equasion,replaced the
pec's and the tube.All power supplies remain constant,the o/p from the
afterglow board remains constant but the o/p from the exponential
amplifier's are moving.The baffling thing about this problem is that the
test signal appears to be okay through the system no drift what so
ever.Between all channels the red seem to be effected the most.I hope you
or one of your colleagues could help me with this problem as ihave asked
nearly everybody i know in europe without much success,also coul you please
let me know how to become a member of the telecine forum.

Regards,
Fergus Mc Carthy
Windmill Lane

--- End of forwarded message from fmcarthy at indigo.ie (fergus mc carthy)




-- 
Rob Lingelbach          |  2660 Hollyridge Dr., Los Angeles, CA 90068
rob at alegria.com  	| "I care not much for a man's religion whose dog or 
rob at sun.alegria.com	|  cat are not the better for it."  --Abraham Lincoln
rob at praia.alegria.com        KB6CUN	   http://www.alegria.com